OpenStudy (anonymous):

Is there any short cut to take out the answer of this->200,000[1+(4/100)]^5=? , without using a calculator,and keeping in mind that there are hardly 30seconds to solve it.

OpenStudy (anonymous):

First: Is this some sort of interest formula you're using? Second: Not really. Type fast.

OpenStudy (anonymous):

200000(1+1/5)=200000+40000=240000 check my math, i did it mentally

OpenStudy (anonymous):

@jm92: u forgot da power'5' sqiggy is there any way you can suggest to work out higher powers fastly??

OpenStudy (anonymous):

and def. manually

OpenStudy (anonymous):

Working out a decimal to the fifth power without using a calculator just takes a lot of time. A suggestion could be to just change the (1+4/100) to (26/25)^5.

OpenStudy (anonymous):

so what about solving it with that big power?

OpenStudy (anonymous):

You can't avoid that.

OpenStudy (anonymous):

ok.thanx :)

OpenStudy (anonymous):

that is the compound interest formula A=P(1+r/n)^(nt) A= total amount P= principal or starting amount r=rate as a decimal n= number of compunding periods per year t=number of years